OpenStudy (anonymous):

if the vectors u = (-2i + 7j) and v=(ai -2j). Determine a: u and v if they are orthogonal u and v is they are parale u and v if the angle between is 2pi/3

OpenStudy (experimentx):

take a dot product ...and fine angle. \[ \vec a \cdot \vec b = |a||b|\cos \theta\]